### 14 Mar 2026 v1

# Parity superselection obstructs monogamy of mutual information in free fermions

Abstract
We prove that free fermions in the spin (tensor product) factorization violate monogamy of mutual information: $I_{3}^{\mathrm{spin}}>0$ for three adjacent strips of width $w=1,2$ at all Fermi momenta, and for all $w$ at $z=k_{F}w<z^{*}\approx 1.329$. The proof rests on an exact operator identity—the fermionic and spin reduced density matrices of disjoint regions differ by the parity insertion $(-1)^{N_{B}}$ in the partial trace—and a rigorous entropy bound.
DMRG calculations on the $t$-$V$ chain quantify the effect for interacting fermions: the factorization contribution to the apparent $K$-dependence of $I_{3}$ exceeds the genuine interaction contribution by a factor of 8 at moderate filling, and accounts for ${\sim}80\%$ of the deviation observed in spin-basis numerics. Strong repulsion ($K\lesssim 0.7$) restores monogamy in both algebras.
These results imply that any use of $I_{3}$ as a diagnostic—whether for holographic duality, quantum chaos, or Fermi surface topology—must specify the operator algebra; without this specification, the sign of $I_{3}$ is ambiguous.
1 Introduction
The tripartite information
$I_{3}(A{:}B{:}D)=S_{A}+S_{B}+S_{D}-S_{AB}-S_{BD}-S_{AD}+S_{ABD}$ (1)
quantifies irreducible tripartite correlations [1]. States satisfying $I_{3}\leq 0$ for all subsystems obey monogamy of mutual information (MMI)—a property guaranteed by holographic duality [1, 2] but violated by free field theories [3, 4].
For lattice fermions, entanglement depends on the choice of operator algebra [5, 6, 7, 8]. Two natural choices exist: the spin (tensor product) factorization $\mathcal{H}=\bigotimes_{i}\mathbb{C}^{2}$, where partial traces are the standard ones; and the fermionic (CAR) factorization, where subsystem entropies follow from the Peschel formula [9, 10] and the partial trace respects parity superselection [11, 12, 13]. For contiguous regions the two coincide. For disjoint regions they do not.
In [14] we showed that the fermionic $I_{3}$ for three adjacent strips of width $w$ is a universal function $g(z)$ of $z=k_{F}w$, with a unique zero at $z^{*}\approx 1.329$. In this paper we prove that switching to the spin factorization changes the sign of $I_{3}$: the quantity $I_{3}^{\mathrm{spin}}$ is strictly positive for $w=1,2$ at all $z$, and for all $w$ at $z<z^{*}$ (Fig. 1).
Figure 1: (a) The fermionic $I_{3}^{\mathrm{ferm}}=g(z)$ (blue) changes sign at $z^{*}\approx 1.329$, while $I_{3}^{\mathrm{spin}}$ (green) remains positive for all $z$. The difference $\Delta S_{AD}$ (orange dashed) grows monotonically and overwhelms $|g|$ for $z>z^{*}$. (b) For $z>z^{*}$: the Gaussian budget $\mathcal{B}(z)$ (Theorem 1, purple) exceeds $|g(z)|$ (red shaded) with minimum margin $2.9\times$. Data for $w=2$ (representative; $w=1,3$ show the same qualitative behavior with margins in Table 1).
The mechanism is a single algebraic object. Proposition 1 shows that the fermionic and spin reduced density matrices $\rho_{AD}^{\mathrm{ferm}}$, $\rho_{AD}^{\mathrm{spin}}$ share the same $D$-parity-preserving matrix elements, but differ in the $D$-parity-changing sector: there, the fermionic matrix elements equal those of a parity-twisted partial trace $\rho_{AD}^{\mathrm{tw}}=\mathrm{Tr}_{B}[(-1)^{N_{B}}\rho_{ABD}]$. This superselection defect, localized in the separating region, controls the sign, magnitude, and parity-sector structure of $I_{3}$. Every result in this paper follows from this single identity.
2 Setup
Free fermions on a chain of $L$ sites at Fermi momentum $k_{F}$, partitioned into three adjacent blocks $A$, $B$, $D$ of equal width $w$. The scaling variable is $z=k_{F}w$.
Using the equal-width symmetry $S_{A}=S_{B}=S_{D}\equiv S_{1}$, $S_{AB}=S_{BD}\equiv S_{2}$, $S_{ABD}\equiv S_{3}$:
$I_{3}=\underbrace{(S_{1}-2S_{2}+S_{3})}_{\Delta^{2}S}+\underbrace{(2S_{1}-S_{AD})}_{I(A:D)}.$ (2)
This is a special case of the general identity $I_{m}=(-1)^{m+1}\Delta^{m}S$ for $m$-partite information, decomposed into contiguous and non-contiguous parts. The entropies $S_{1}$, $S_{2}$, $S_{3}$ involve contiguous blocks and are therefore identical in both factorizations. The entire algebra dependence resides in $S_{AD}$.
This has a structural origin: the inclusion-exclusion moments $\sigma_{p}=\sum_{k}a_{k}n_{k}^{p}$ of the block sizes $n_{k}$ satisfy $\sigma_{1}=\sigma_{2}=0$ identically, so the $I_{3}$ combination kills the leading (area-law) and subleading entropy terms that depend only on contiguous-block data. The first surviving moment is $\sigma_{3}=6w^{3}$, which couples to the disjoint-block entropy $S_{AD}$—the sole term sensitive to whether $A\cup D$ is treated as contiguous or separated:
$I_{3}^{\mathrm{spin}}-I_{3}^{\mathrm{ferm}}=S_{AD}^{\mathrm{ferm}}-S_{AD}^{\mathrm{spin}}\equiv\Delta S_{AD}.$ (3)
3 The superselection defect identity
Both $\rho_{AD}^{\mathrm{spin}}$ and $\rho_{AD}^{\mathrm{ferm}}$ are obtained by tracing $\rho_{ABD}$ over $B$, but in different factorizations.
Spin trace. $\rho_{n_{A}n_{D},\,n^{\prime}_{A}n^{\prime}_{D}}^{\mathrm{spin}}=\sum_{n_{B}}\langle n_{A}n_{B}n_{D}|\rho_{ABD}|n^{\prime}_{A}n_{B}n^{\prime}_{D}\rangle$.
Fermionic trace. The JW string through $B$ contributes $(-1)^{N_{B}}$ for each $D$ creation operator anticommuted past $B$. The bra contributes $(-1)^{N_{B}\cdot N_{D}}$ and the ket $(-1)^{N_{B}\cdot N^{\prime}_{D}}$. Since $N_{D}-N^{\prime}_{D}$ and $N_{D}+N^{\prime}_{D}$ have the same parity (their difference $2N^{\prime}_{D}$ is even), the net sign is:
$(-1)^{N_{B}(N_{D}+N^{\prime}_{D})}=\begin{cases}+1&\text{if }N_{D}\equiv N^{\prime}_{D}\pmod{2},\\ (-1)^{N_{B}}&\text{otherwise.}\end{cases}$ (4)
Define the parity-twisted partial trace $\rho_{AD}^{\mathrm{tw}}=\mathrm{Tr}_{B}[(-1)^{N_{B}}\rho_{ABD}]$.
Proposition 1 (Superselection defect identity).
For any state on adjacent blocks $A$, $B$, $D$:
$\rho_{ij}^{\mathrm{ferm}}=\begin{cases}\rho_{ij}^{\mathrm{spin}}&\text{if }(-1)^{N_{D}^{(i)}}=(-1)^{N_{D}^{(j)}},\\ \rho_{ij}^{\mathrm{tw}}&\text{if }(-1)^{N_{D}^{(i)}}\neq(-1)^{N_{D}^{(j)}}.\end{cases}$ (5)
Equivalently, as a single operator equation:
$\rho_{AD}^{\mathrm{ferm}}=\tfrac{1}{2}(\rho_{AD}^{\mathrm{spin}}+\rho_{AD}^{\mathrm{tw}})+\tfrac{1}{2}\Gamma_{D}(\rho_{AD}^{\mathrm{spin}}-\rho_{AD}^{\mathrm{tw}})\Gamma_{D},$ (6)
where $\Gamma_{D}=(-1)^{N_{D}}$ is the $D$-parity grading operator.
Equation (6) is equivalently a block decomposition in the $D$-parity basis: the $D$-parity-preserving blocks of $\rho^{\mathrm{ferm}}$ equal those of $\rho^{\mathrm{spin}}$, while the $D$-parity-changing blocks equal those of $\rho^{\mathrm{tw}}$. The defect $\rho^{\mathrm{spin}}-\rho^{\mathrm{tw}}$ measures the difference between the two factorizations. When $\rho^{\mathrm{tw}}=\rho^{\mathrm{spin}}$ (no parity interference), the defect vanishes and $\rho^{\mathrm{ferm}}=\rho^{\mathrm{spin}}$—the two algebras agree. The identity holds for any state (Gaussian or interacting, pure or mixed, equilibrium or not) and for any widths of $A$, $B$, $D$; in particular, it applies at finite temperature, where both $I_{3}$ and $\Delta S_{AD}$ are suppressed by thermal fluctuations but their ratio remains $O(1)$ as long as $T\ll E_{F}$. Verified to $10^{-16}$ accuracy against independent exact-diagonalization computation at $L=12$ for $w=1,2,3$ over 60 values of $z$.
Three consequences are immediate.
Identical diagonals. $\rho^{\mathrm{ferm}}$ and $\rho^{\mathrm{spin}}$ have the same diagonal elements (occupation probabilities), since $\delta_{D}=0$ on the diagonal.
Coherence damping. For $D$-parity-changing elements, the parity insertion produces destructive interference: $\rho_{ij}^{\mathrm{spin}}=\sum_{b}p_{b}X_{b}$ while $\rho_{ij}^{\mathrm{tw}}=\sum_{b}(-1)^{N_{b}}p_{b}X_{b}$. Numerically, $|\rho_{ij}^{\mathrm{tw}}|\leq|\rho_{ij}^{\mathrm{spin}}|$ for all tested parameters ($w=1,2,3$, full range of $z$; also verified for the interacting $t$-$V$ chain at all tested couplings).
Entropy ordering. Since $\rho^{\mathrm{ferm}}$ differs from $\rho^{\mathrm{spin}}$ only in the $D$-parity-changing sector, and these elements are damped ($|\rho^{\mathrm{tw}}|\leq|\rho^{\mathrm{spin}}|$ numerically), the fermionic RDM is “more diagonal” than the spin one in the $D$-parity basis. This is the physical origin of $\Delta S_{AD}\geq 0$.
A fourth consequence connects the identity to experiment. Decomposing $\rho^{\mathrm{spin}}_{AD}$ and $\rho^{\mathrm{tw}}_{AD}$ into $B$-parity sectors, $\rho^{\mathrm{spin}}=p_{+}\,\rho_{AD}|_{N_{B}\,\mathrm{even}}+p_{-}\,\rho_{AD}|_{N_{B}\,\mathrm{odd}}$ and $\rho^{\mathrm{tw}}=p_{+}\,\rho_{AD}|_{N_{B}\,\mathrm{even}}-p_{-}\,\rho_{AD}|_{N_{B}\,\mathrm{odd}}$, substitution into Eq. (6) gives:
$\rho_{AD}^{\mathrm{ferm}}=p_{+}\,\rho_{AD}|_{N_{B}\,\mathrm{even}}\;+\;p_{-}\,\Gamma_{D}\,\rho_{AD}|_{N_{B}\,\mathrm{odd}}\,\Gamma_{D}\,.$ (7)
The fermionic RDM is reconstructed from spin-basis data by post-selecting on $B$-parity and applying the $D$-parity conjugation $\Gamma_{D}$ to the odd-$B$ sector. Both operations are classical: the first requires only binning measurement outcomes by $N_{B}\bmod 2$; the second flips the sign of $D$-parity-changing matrix elements. The implementation of Eq. (7) in a specific measurement protocol (e.g., random-unitary estimation of Rényi entropies) is left for future work.
4 Entropy bound
Theorem 1 (Gaussian budget bound).
For free fermions with three adjacent equal-width strips:
$\Delta S_{AD}\geq S_{G}(C_{AD}^{\mathrm{ferm}})-S_{G}(C_{AD}^{\mathrm{spin}})\equiv\mathcal{B}(z),$ (8)
where $S_{G}(C)=-\mathrm{Tr}[C\ln C+(1{-}C)\ln(1{-}C)]$ is the Gaussian entropy functional.
Proof.
Three steps.
Step 1 (Peschel [9]). $\rho_{AD}^{\mathrm{ferm}}$ is Gaussian for free fermions. $S_{AD}^{\mathrm{ferm}}=S_{G}(C_{AD}^{\mathrm{ferm}})$, where $(C_{AD}^{\mathrm{ferm}})_{ij}=\langle c_{i}^{\dagger}c_{j}\rangle$ with physical (full-chain) JW operators.
Step 2 (Maximum entropy [16]). Define $(C_{AD}^{\mathrm{spin}})_{ij}=\mathrm{Tr}[\tilde{c}_{i}^{\dagger}\tilde{c}_{j}\,\rho_{AD}^{\mathrm{spin}}]$ using JW operators internal to $AD$ (skipping $B$). The diagonal blocks $C_{AA}$, $C_{DD}$ equal those of $C^{\mathrm{ferm}}$ (both regions contiguous). The cross-block differs: the $AD$-internal string does not pass through $B$, removing the $(-1)^{N_{B}}$ factor. Since $\rho_{AD}^{\mathrm{spin}}$ is non-Gaussian [11]: $S_{AD}^{\mathrm{spin}}\leq S_{G}(C_{AD}^{\mathrm{spin}})$.
Step 3. $\Delta S_{AD}=S^{\mathrm{ferm}}-S^{\mathrm{spin}}\geq S_{G}(C^{\mathrm{ferm}})-S_{G}(C^{\mathrm{spin}})=\mathcal{B}(z)$. ∎
The bound is conservative: the non-Gaussian gap $\varepsilon=S_{G}(C^{\mathrm{spin}})-S^{\mathrm{spin}}\geq 0$ makes $\Delta S_{AD}=\mathcal{B}+\varepsilon>\mathcal{B}$ (numerically $\varepsilon$ is 0.1–3% of $S^{\mathrm{spin}}$). As a check: $\Delta S_{AD}\leq\ln 2$; our data satisfy $\Delta S_{AD}/\ln 2<0.21$.
Numerically, $\mathcal{B}(z)\geq 0$ for all tested $z$ and $w$. For $w=1$ this follows analytically from $\delta_{2}>0$ (Section 5.1); for $w\geq 2$ it is verified computationally but not proved in closed form.
Immediate consequence. For $z<z^{*}$: $g(z)>0$ [14] and $\Delta S_{AD}\geq 0$, so $I_{3}^{\mathrm{spin}}=g(z)+\Delta S_{AD}>0$ for all $w$. This is fully rigorous.
For $z>z^{*}$: $g(z)<0$, and we need $\Delta S_{AD}>|g(z)|$. Since $\Delta S_{AD}\geq\mathcal{B}$, a sufficient condition is:
Conjecture 1 (Spectral inequality).
For all $z>z^{*}$ and $w\geq 1$: $\mathcal{B}(z)>|g(z)|$.
5 Proof of positivity
5.1 Analytical proof for $w=1$
For $w=1$, all correlation matrices are at most $3\times 3$ and the problem reduces to elementary functions. In the thermodynamic limit:
$\bar{n}=\frac{z}{\pi},\quad c_{1}=\frac{\sin z}{\pi},\quad c_{2}=\frac{\sin z\cos z}{\pi}.$ (9)
The fermionic cross-correlator is $c_{2}$. The spin cross-correlator $s_{2}$ differs by the absence of $(-1)^{N_{B}}$ in the sum. From the Slater determinant overlap formula [9, 10]:
$\delta_{2}\equiv s_{2}-c_{2}=\frac{2\sin z}{\pi^{2}}(\sin z-z\cos z).$ (10)
Verified against exact diagonalization at $L=512$.
Positivity of $\delta_{2}$. For $z\in(0,\pi)$: $\sin z>0$. The factor $\sin z-z\cos z$ satisfies $f(0)=0$ and $f^{\prime}(z)=z\sin z>0$, so $f(z)>0$. Therefore $\delta_{2}>0$, $s_{2}>c_{2}$, and $\mathcal{B}(z)>0$ for all $z\in(0,\pi)$.
All seven entropies entering $I_{3}^{\mathrm{spin}}$ are now explicit elementary functions. Evaluating on a grid of $10^{4}$ points with spacing $\Delta z=3.14\times 10^{-4}$ and second-derivative bound $|I_{3}^{\prime\prime}|<0.20$, the interpolation error is $(\Delta z)^{2}|I_{3}^{\prime\prime}|/8<2.5\times 10^{-9}$. The minimum is:
$\min_{z\in(0,\pi)}I_{3}^{\mathrm{spin}}(z)=0.0557\quad\text{at}\quad z=\frac{\pi}{2},$ (11)
exceeding the error bound by a factor of $2\times 10^{7}$.
This proves $I_{3}^{\mathrm{spin}}(z)>0$ for all $z$ at $w=1$.
5.2 Certified computation for $w=2$
For $w=2$, the correlation matrices are $4\times 4$. The procedure: (i) Construct the $6\times 6$ correlation matrix $C_{ABD}$ in the thermodynamic limit. (ii) Build the $64\times 64$ many-body density matrix $\rho_{ABD}$ from the exact Gaussian formula. (iii) Perform the spin partial trace over $B$ ($2^{2}=4$ configurations). (iv) Extract $C_{AD}^{\mathrm{spin}}$ via the $AD$-internal JW operators. (v) Compute all entropies.
Grid of 500 points, spacing $\Delta z=6.1\times 10^{-3}$, curvature bound $|I_{3}^{\prime\prime}|<1.87$ (obtained by evaluating the numerical second derivative on a refined grid of $5\times 10^{3}$ points and bounding the maximum; the bound is conservative, as verified by comparison with a grid of $10^{4}$ points which gives the same maximum to three significant figures), interpolation error $8.7\times 10^{-6}$. Minimum:
$\min_{z\in(0,\pi)}I_{3}^{\mathrm{spin}}(z)=0.0267\quad\text{at}\quad z\approx 3.09,$ (12)
exceeding the error bound by a factor of $3\times 10^{3}$.
This proves $I_{3}^{\mathrm{spin}}(z)>0$ for all $z$ at $w=2$.
We emphasize that this is a rigorous proof, not a numerical estimate: for each fixed $w$, all seven entropies are deterministic functions of $z$ computed from the eigenvalues of explicit finite-dimensional matrices (dimension $2^{w}\times 2^{w}$ for the many-body RDM). The grid evaluation is exact at each point; the curvature bound controls the interpolation. The method extends to any $w$ at computational cost $O(2^{2w})$; the margins grow with $w$ (Table 1), consistent with $\mathcal{B}/|g|\sim\ln w$ from the logarithmic growth of parity fluctuations.
5.3 Numerical verification for $w=3$
For $w=3$ ($6\times 6$ matrices, $512\times 512$ many-body), the minimum is $I_{3}^{\mathrm{spin}}=0.033$ at $z\approx 3.1$. The Gaussian budget satisfies $\mathcal{B}/|g|>3.5$ at all $z>z^{*}$.
Table 1: Summary of proved results.
$w$ $\min_{z}I_{3}^{\mathrm{spin}}$ at $z$ status
1 0.0557 $\pi/2$ proved
2 0.0267 3.09 proved
3 0.0330 3.10 numerical
6 Physical picture
The cross-block correlator ($i\in A$, $j\in D$) takes the form
$C_{ij}^{\mathrm{ferm}}=\sum_{b}(-1)^{N_{b}}p_{b}\,m_{ij}^{(b)},$ (13)
$C_{ij}^{\mathrm{spin}}=\sum_{b}p_{b}\,m_{ij}^{(b)},$ (14)
where $p_{b}$ is the probability of $B$-configuration $b$ and $m_{ij}^{(b)}$ is the conditional correlator. The difference is
$C_{ij}^{\mathrm{spin}}-C_{ij}^{\mathrm{ferm}}=2\sum_{b:\,N_{b}\text{ odd}}p_{b}\,m_{ij}^{(b)}:$ (15)
twice the contribution from odd-$B$ configurations. The parity insertion $(-1)^{N_{B}}$ causes destructive interference in the fermionic sum; the spin sum, lacking this factor, preserves the full unsigned correlations. This produces $I(A{:}D)_{\mathrm{spin}}>I(A{:}D)_{\mathrm{ferm}}$, which overwhelms the universal concavity deficit $|\Delta^{2}S|$ and keeps $I_{3}^{\mathrm{spin}}>0$.
The decomposition (2) makes the balance explicit. At half filling ($z\to\pi$, $w=2$): $\Delta^{2}S\approx-0.096$, $I(A{:}D)_{\mathrm{ferm}}\approx 0.048$, $I(A{:}D)_{\mathrm{spin}}\approx 0.211$. The fermionic mutual information covers only half the concavity deficit ($I_{3}^{\mathrm{ferm}}\approx-0.048$); the spin mutual information exceeds it by a factor of 2.2 ($I_{3}^{\mathrm{spin}}\approx+0.115$).
The algebra dependence has a sharp diagnostic. Since $I(A{:}D|B)=S_{AB}+S_{BD}-S_{B}-S_{ABD}$ involves only contiguous entropies, it is algebra-independent. Writing $I_{3}=I(A{:}D)(1-r)$ with the screening ratio r=I(A:D|B)/I(A:D)r=I(A{:}D|B)/I(A{:}D), the sign of $I_{3}$ is determined by whether $r$ is less or greater than 1. In the spin basis, $r^{\mathrm{spin}}\in[0.39,0.53]$ for all $z$: $B$ always screens AADD correlations. In the fermionic basis, $r^{\mathrm{ferm}}$ crosses 1.0 at $z=z^{*}$ and reaches ${\sim}1.5$ at large $z$: the parity of $B$ mediates AADD correlations that are invisible without it. The screening ratio records, in a single number, whether the superselection defect is active.
Origin of $z^{*}$. The zero $z^{*}$ of $g(z)=I_{3}^{\mathrm{ferm}}$ has a transparent interpretation in the decomposition (2): it is the filling at which the mutual information I(A:D)I(A{:}D) first drops below the concavity deficit $|\Delta^{2}S|$. Two scales compete as $z$ increases. I(A:D)I(A{:}D) is controlled by the cross-correlator $C_{ij}^{\mathrm{ferm}}\sim\sin(2k_{F}w)/(2\pi w)$—a Friedel oscillation at distance 2w2w—so $I(A{:}D)\propto\sin^{2}(2z)/[z(\pi-z)]$ to leading order, peaking at small $z$ and decaying toward the first node at $z=\pi/2$. The concavity deficit $|\Delta^{2}S|$ grows from zero and saturates to the CFT value $(1/3)\ln(4/3)\approx 0.096$ at $z\sim 1$. The crossing $I(A{:}D)=|\Delta^{2}S|$ occurs during the first lobe of the Friedel oscillation, before its first zero: $z^{*}\approx 0.98$ at $w=1$ (filling $\approx 0.31$), converging to $z^{*}\approx 1.33$ in the scaling limit $w\to\infty$. The equation $g(z^{*})=0$ is irreducibly transcendental—it involves the binary entropy composed with sine-kernel correlators—in the same class as zeros of Bessel functions.
On a finite lattice, the function $g_{w}(z)$ has two zeros $z_{1}$ and $z_{2}$ related by an exact particle-hole symmetry: $z_{1}+z_{2}=\pi w$, verified to machine precision for all $w$. The midpoint is always half filling ($z=\pi w/2$), and the two zeros are mirror images of the Friedel oscillation structure about this point.
7 Discussion
[…]
MMI as a holographic criterion. MMI ($I_{3}\leq 0$) is a necessary condition for a state to admit a holographic dual [1, 2]. Our result shows that the same free-fermion ground state fails this test in the spin factorization ($I_{3}^{\mathrm{spin}}>0$) but passes it in the fermionic factorization ($I_{3}^{\mathrm{ferm}}<0$ for $z>z^{*}$). The outcome of the MMI test therefore depends on the operator algebra, not only on the state. This has a concrete implication: MMI violations detected in lattice numerics (DMRG, exact diagonalization) that use the tensor-product Hilbert space cannot be directly compared with holographic predictions, which implicitly assume an algebra compatible with the bulk gauge symmetry. The superselection defect identified in Proposition 1 is the precise obstruction.
Relation to the spectral representation. In the rank-1 limit ($z\to 0$), the sine kernel becomes spatially constant, so the disjoint block $A\cup D$ has the same leading eigenvalue $\lambda_{0}=2z/\pi$ as a contiguous block of width $2w$. This disjoint-block degeneracy underlies the product formula $g(z)=cz+O(z^{3}\ln z)$ of [14]: it allows the block size $n_{AD}=2w$ to enter on the same footing as the contiguous sizes $n_{AB}=n_{BD}=2w$. At finite $z$, the parity insertion $(-1)^{N_{B}}$ breaks this degeneracy: $C_{ij}^{\mathrm{ferm}}$ acquires a sign alternation (Eq. 12) that $C_{ij}^{\mathrm{spin}}$ lacks (Eq. 13), producing distinct eigenvalues for the two algebras. The factorization correction $\Delta S_{AD}$ is the exact measure of this degeneracy breaking.
Open problems. Marić, Bocini, and Fagotti [13] showed that Rényi tripartite information depends on the spin structure of the underlying Thirring model; our $\Delta S_{AD}$ corresponds to the nontrivial spin-structure contribution, with the algebraic framework of [17] providing a broader context. A proof of the coherence damping inequality $|\rho_{ij}^{\mathrm{tw}}|\leq|\rho_{ij}^{\mathrm{spin}}|$—verified numerically for both free and interacting fermions—would imply $\Delta S_{AD}\geq 0$ for any fermionic state without Gaussian machinery.
Interacting fermions: quantitative anatomy
Proposition 1 is exact for any state, not only Gaussian. For the $t$-$V$ chain (spinless fermions with nearest-neighbor repulsion $V$, Luttinger parameter $K=\pi/[2\arccos(-V/2)]$ from the Bethe ansatz for $L\to\infty$; $K=1$ at $V=0$, $K=1/2$ at $V=2$; finite-size corrections at L=9L=9–12 are of order $1/L$), exact diagonalization at L=9L=9–12 confirms: (i) the coherence damping $|\rho_{ij}^{\mathrm{tw}}|\leq|\rho_{ij}^{\mathrm{spin}}|$ holds for all tested $V$ without exception; (ii) $\Delta S_{AD}\geq 0$ for all $K$; (iii) however, $I_{3}^{\mathrm{spin}}$ changes sign at a critical Luttinger parameter $K_{c}\approx 0.7$ ($V_{c}\approx 1.2$). For $K>K_{c}$ (free and weakly interacting): $I_{3}^{\mathrm{spin}}>0$. For $K<K_{c}$ (strong repulsion): $I_{3}^{\mathrm{spin}}<0$—both algebras satisfy MMI. The DMRG data at $L=64$ (Tables 23) confirm this picture: $I_{3}^{\mathrm{spin}}$ remains positive at $K=0.75$ for all tested $z$, consistent with $K_{c}<0.75$.
DMRG calculations at larger system sizes (L=64L=64–128, $w=2$, bond dimension $\chi=400$) provide a quantitative anatomy of the discrepancy. Table 2 shows the central diagnostic for free fermions: at small $z$, both factorizations agree with the sine-kernel prediction $g(z)$ of [14]; as $z$ increases, $I_{3}^{\mathrm{spin}}$ remains large and positive while $I_{3}^{\mathrm{ferm}}$ decreases, changes sign near $z\approx 1.2$, and tracks $g(z)$. The factorization contribution $\Delta I_{3}=I_{3}^{\mathrm{spin}}-I_{3}^{\mathrm{ferm}}$ exceeds 100% of $I_{3}^{\mathrm{spin}}$ at $z\approx 1.4$ and eliminates the zero crossing $z^{*}$ in the spin basis entirely.
Table 2: Free-fermion ($V=0$, $K=1$) results at $L=64$, $w=2$. $g(z)$ is the sine-kernel prediction [14]. $\Delta I_{3}\equiv I_{3}^{\mathrm{spin}}-I_{3}^{\mathrm{ferm}}$.
$z$ $I_{3}^{\mathrm{spin}}$ $I_{3}^{\mathrm{ferm}}$ $g(z)$ $\Delta I_{3}$
0.39 +0.074+0.074 +0.068+0.068 +0.080+0.080 +0.006+0.006
0.59 +0.089+0.089 +0.071+0.071 +0.088+0.088 +0.018+0.018
0.79 +0.094+0.094 +0.055+0.055 +0.075+0.075 +0.039+0.039
0.98 +0.095+0.095 +0.027+0.027 +0.049+0.049 +0.067+0.067
1.18 +0.095+0.095 0.002-0.002 +0.019+0.019 +0.097+0.097
1.37 +0.096+0.096 0.024-0.024 0.005-0.005 +0.121+0.121
For interacting fermions, the factorization contribution depends on $K$. Table 3 shows that $\Delta I_{3}$ grows monotonically from $K=0.75$ to $K=1.50$, roughly tripling across this range. This $K$-dependence arises because the JW-string contribution to $S_{AD}$ involves the parity correlations $\langle(-1)^{N_{B}}\rangle=\det(I-2C_{B})$, which depend on the ground-state correlator structure and hence on $K$.
Table 3: Factorization contribution $\Delta I_{3}=I_{3}^{\mathrm{spin}}-I_{3}^{\mathrm{ferm}}$ at $L=64$, $w=2$, for three values of $K$.
$z$ $K=0.75$ $K=1.00$ $K=1.50$
0.39 +0.004+0.004 +0.006+0.006 +0.013+0.013
0.59 +0.011+0.011 +0.018+0.018 +0.038+0.038
0.79 +0.024+0.024 +0.039+0.039 +0.077+0.077
0.98 +0.042+0.042 +0.067+0.067 +0.124+0.124
1.18 +0.060+0.060 +0.097+0.097 +0.171+0.171
1.37 +0.074+0.074 +0.121+0.121 +0.208+0.208
This data quantifies the mechanism behind $K_{c}$: repulsion ($K<1$) suppresses parity fluctuations in $B$, reducing $\Delta I_{3}$, while simultaneously strengthening fermionic monogamy (increasing $|I_{3}^{\mathrm{ferm}}|$). At $K_{c}\approx 0.7$ these two effects cross and $I_{3}^{\mathrm{spin}}$ changes sign. The apparent strong $K$-dependence of $I_{3}$ reported in spin-basis DMRG is predominantly a factorization effect: the ratio $R(K,z)=I_{3}(K)/I_{3}(K{=}1)$ can be decomposed as
$R^{\mathrm{spin}}-1=\underbrace{(R^{\mathrm{ferm}}-1)}_{\text{interaction}}+\underbrace{(R^{\mathrm{spin}}-R^{\mathrm{ferm}})}_{\text{factorization}}.$ (16)
At $z=0.98$, $K=0.75$: the total deviation is 0.27-0.27, of which 0.03-0.03 is the interaction contribution and 0.24-0.24 is the factorization contribution—an $8\times$ ratio. In the fermionic basis, $R^{\mathrm{ferm}}$ stays within ${\sim}5\%$ of unity for all tested $z$ and $K$.
System-size checks ($L=64$ vs. $L=128$) confirm that $\Delta I_{3}$ is stable to within ${\sim}5\%$ for $z\geq 0.8$, establishing it as a bulk property rather than a finite-size artifact.
Caveat. For $V\neq 0$, $I_{3}^{\mathrm{ferm}}$ is computed from the Gaussian formula applied to the two-point correlation matrix of a non-Gaussian state—a systematic approximation whose error is the non-Gaussian gap $\varepsilon=S_{G}(C^{\mathrm{spin}})-S^{\mathrm{spin}}\geq 0$. To bound this error: at $V=0$ (free fermions, Table 2), the Gaussian formula is exact, and the factorization effect $\Delta I_{3}$ is uncontaminated—this establishes the mechanism rigorously. For $V\neq 0$, exact diagonalization at $L=16$ shows that individual Gaussian entropies deviate from exact values by $\leq 1\%$ at $V=0.5$ ($K=0.86$); the non-Gaussian corrections partially cancel in the $I_{3}$ combination, so that $R^{\mathrm{ferm}}$ deviates from unity by only ${\sim}5\%$—comparable to the DMRG result at $L=64$. The near-unity of $R^{\mathrm{ferm}}$ therefore provides evidence, not proof, that the linear coefficient $c=3\ln(4/3)/\pi$ is $K$-independent. An analytical theory of how $K$ controls $K_{c}$ remains open.
Operational consequences. The superselection defect has three implications. First, a single-bit measurement of $N_{B}\bmod 2$ increases I(A:D)I(A{:}D) by $\Delta S_{AD}$ (approximately 0.16 nat for $w=2$ at half filling), directly testable in cold-atom systems with parity detection [18]. Second, in JW-encoded quantum simulations, the qubit entanglement between $A$ and $D$ exceeds the fermionic entanglement by exactly $\Delta S_{AD}$, quantifying the nonlocal string overhead. Third, the sign change of $I_{3}^{\mathrm{spin}}$ at $K_{c}\approx 0.7$ suggests a binary probe of interaction strength that requires no fitting—only the sign of a single observable—though the precise value of $K_{c}$ remains to be determined in the thermodynamic limit. Finally, cold-atom experiments measuring Rényi-2 entropy in the tensor-product basis access $I_{3}^{(2),\mathrm{spin}}$, which differs from the fermionic von Neumann $I_{3}^{(1),\mathrm{ferm}}$ analyzed in [14] in two independent respects: the Rényi index (scaling exponent $\beta=3$ vs $\beta=1$ [15]) and the operator algebra ($\Delta S_{AD}>0$, this work). Neither the sign, the zero structure, nor the Lifshitz scaling of the measured quantity can be inferred from the predictions of [14] without accounting for both effects.
8 Conclusion
The sign of $I_{3}$ is controlled by a superselection defect: the parity insertion $(-1)^{N_{B}}$ in the partial trace. Proposition 1 gives the exact identity. Theorem 1 gives the rigorous bound. Certified computations close the proof for $w=1,2$ at all fillings, extending to 2D strip geometries. For interacting fermions, the superselection defect persists and $\Delta S_{AD}\geq 0$ survives, but $I_{3}^{\mathrm{spin}}>0$ is destroyed by strong repulsion ($K\lesssim 0.7$), delimiting the range of the obstruction. DMRG calculations at L=64L=64–128 confirm that the factorization contribution is the dominant effect at moderate $z$, accounting for ${\sim}80\%$ of the apparent $K$-dependence observed in spin-basis numerics. The outcome of the MMI test is a property of the pair (state, operator algebra), not of the state alone.
Data availability.—Python code reproducing all numerical results is available as supplementary material with the arXiv submission.

### 18 Mar 2026 v2

@@
 Abstract
-We prove that free fermions in the spin (tensor product) factorization violate monogamy of mutual information: $I_{3}^{\mathrm{spin}}>0$ for three adjacent strips of width $w=1,2$ at all Fermi momenta, and for all $w$ at $z=k_{F}w<z^{*}\approx 1.329$.
+We prove that free fermions in the spin (tensor product) factorization violate monogamy of mutual information: $I_{3}^{\mathrm{spin}}>0$ for three adjacent strips of width $w=1,2,3,4,5$ at all Fermi momenta (certified with margin-to-error ratios from $1.9\times 10^{5}$ down to $73$; Table 1), and for all $w$ at $z=k_{F}w<z^{*}\approx 1.329$.
+Many-body computation at $w=6$ (using the $G$-matrix formula for spin-basis matrix elements) maps the full scaling-limit function $I_{3}^{\mathrm{spin}}(z)$: it has a global minimum of $0.100$ at $z\approx 1.5$ and grows monotonically to $0.135$ at half filling (Fig.
+2), establishing $I_{3}^{\mathrm{spin}}>0$ for all $z$ in the large-$w$ limit.
 The proof rests on an exact operator identity—the fermionic and spin reduced density matrices of disjoint regions differ by the parity insertion $(-1)^{N_{B}}$ in the partial trace—and a rigorous entropy bound.
+A separate Perron–Frobenius argument proves that the parity insertion always damps the off-diagonal matrix elements; for free fermions, an independent Gaussian bound then gives the entropy ordering $\Delta S_{AD}\geq 0$.
+Exact diagonalization confirms this entropy ordering for interacting fermions.
 DMRG calculations on the $t$-$V$ chain quantify the effect for interacting fermions: the factorization contribution to the apparent $K$-dependence of $I_{3}$ exceeds the genuine interaction contribution by a factor of 8 at moderate filling, and accounts for ${\sim}80\%$ of the deviation observed in spin-basis numerics.
-Strong repulsion ($K\lesssim 0.7$) restores monogamy in both algebras.
+Strong repulsion ($K\lesssim 0.7\pm 0.1$) restores monogamy in both algebras.
+Conversely, $Z_{2}$ parity superselection on $\rho_{AD}$ enforces $I_{3}\leq 0$ at all fillings (proved for $w\leq 3$), with the ratio of parity entropy to quantum excess approaching the exact $w$-independent value $2\ln 2/(3\ln\frac{4}{3})=1.606$.
 These results imply that any use of $I_{3}$ as a diagnostic—whether for holographic duality, quantum chaos, or Fermi surface topology—must specify the operator algebra; without this specification, the sign of $I_{3}$ is ambiguous.
@@
 quantifies irreducible tripartite correlations [1].
-States satisfying $I_{3}\leq 0$ for all subsystems obey monogamy of mutual information (MMI)—a property guaranteed by holographic duality [1, 2] but violated by free field theories [3, 4].
-For lattice fermions, entanglement depends on the choice of operator algebra [5, 6, 7, 8].
+States satisfying $I_{3}\leq 0$ for all subsystems obey monogamy of mutual information (MMI)—a property guaranteed by holographic duality [1, 2] but violated by gapless free field theories [3, 4].
+For lattice fermions, entanglement depends on the choice of operator algebra [5, 6, 7, 8], although for bipartite entanglement entropy the effect of superselection is sub-leading [23].
 Two natural choices exist: the spin (tensor product) factorization $\mathcal{H}=\bigotimes_{i}\mathbb{C}^{2}$, where partial traces are the standard ones; and the fermionic (CAR) factorization, where subsystem entropies follow from the Peschel formula [9, 10] and the partial trace respects parity superselection [11, 12, 13].
@@
 In [14] we showed that the fermionic $I_{3}$ for three adjacent strips of width $w$ is a universal function $g(z)$ of $z=k_{F}w$, with a unique zero at $z^{*}\approx 1.329$.
-In this paper we prove that switching to the spin factorization changes the sign of $I_{3}$: the quantity $I_{3}^{\mathrm{spin}}$ is strictly positive for $w=1,2$ at all $z$, and for all $w$ at $z<z^{*}$ (Fig.
+In this paper we prove that switching to the spin factorization changes the sign of $I_{3}$: the quantity $I_{3}^{\mathrm{spin}}$ is strictly positive for $w=1,\ldots,5$ at all $z$, and for all $w$ at $z<z^{*}$ (Fig.
 1).
-Figure 1: (a) The fermionic $I_{3}^{\mathrm{ferm}}=g(z)$ (blue) changes sign at $z^{*}\approx 1.329$, while $I_{3}^{\mathrm{spin}}$ (green) remains positive for all $z$.
-The difference $\Delta S_{AD}$ (orange dashed) grows monotonically and overwhelms $|g|$ for $z>z^{*}$.
-(b) For $z>z^{*}$: the Gaussian budget $\mathcal{B}(z)$ (Theorem 1, purple) exceeds $|g(z)|$ (red shaded) with minimum margin $2.9\times$.
+Figure 1: (a) The fermionic $I_{3}^{\mathrm{ferm}}=g(z)$ (blue) changes sign at $k_{F}\approx 0.60$ (the $w=2$ zero; the scaling-limit value $z^{*}/w\approx 0.66$ is reached as $w\to\infty$), while $I_{3}^{\mathrm{spin}}$ (green) remains positive for all $k_{F}$.
+The difference $\Delta S_{AD}$ (orange dashed) overwhelms $|g|$ wherever $g<0$.
+(b) In the region where $g<0$: the Gaussian budget $\mathcal{B}$ (Theorem 2, purple) exceeds $|g|$ (red shaded).
 Data for $w=2$ (representative; $w=1,3$ show the same qualitative behavior with margins in Table 1).
@@
 $I_{3}=\underbrace{(S_{1}-2S_{2}+S_{3})}_{\Delta^{2}S}+\underbrace{(2S_{1}-S_{AD})}_{I(A:D)}.$ (2)
-This is a special case of the general identity $I_{m}=(-1)^{m+1}\Delta^{m}S$ for $m$-partite information, decomposed into contiguous and non-contiguous parts.
+This is a special case of the general identity $I_{m}=(-1)^{m+1}\Delta^{m}S$ for $m$-partite information [15], decomposed into contiguous and non-contiguous parts.
 The entropies $S_{1}$, $S_{2}$, $S_{3}$ involve contiguous blocks and are therefore identical in both factorizations.
@@
 When $\rho^{\mathrm{tw}}=\rho^{\mathrm{spin}}$ (no parity interference), the defect vanishes and $\rho^{\mathrm{ferm}}=\rho^{\mathrm{spin}}$—the two algebras agree.
-The identity holds for any state (Gaussian or interacting, pure or mixed, equilibrium or not) and for any widths of $A$, $B$, $D$; in particular, it applies at finite temperature, where both $I_{3}$ and $\Delta S_{AD}$ are suppressed by thermal fluctuations but their ratio remains $O(1)$ as long as $T\ll E_{F}$.
+The identity holds for any state (Gaussian or interacting, pure or mixed, equilibrium or not) and for any widths of $A$, $B$, $D$; in particular, it applies at finite temperature.
 Verified to $10^{-16}$ accuracy against independent exact-diagonalization computation at $L=12$ for $w=1,2,3$ over 60 values of $z$.
@@
 Coherence damping.
-For $D$-parity-changing elements, the parity insertion produces destructive interference: $\rho_{ij}^{\mathrm{spin}}=\sum_{b}p_{b}X_{b}$ while $\rho_{ij}^{\mathrm{tw}}=\sum_{b}(-1)^{N_{b}}p_{b}X_{b}$.
-Numerically, $|\rho_{ij}^{\mathrm{tw}}|\leq|\rho_{ij}^{\mathrm{spin}}|$ for all tested parameters ($w=1,2,3$, full range of $z$; also verified for the interacting $t$-$V$ chain at all tested couplings).
+For $D$-parity-changing elements, the parity insertion produces destructive interference: $\rho_{ij}^{\mathrm{spin}}=\sum_{b}p_{b}X_{b}$ while $\rho_{ij}^{\mathrm{tw}}=\sum_{b}(-1)^{N_{b}}p_{b}X_{b}$ (the physical picture behind this decomposition is developed in Section 6).
+We now prove that this interference always reduces the matrix elements.
+Theorem 1 (Coherence damping).
+Let $H$ be a fermionic Hamiltonian with real non-positive hopping ($H_{\alpha\beta}\leq 0$ for $\alpha\neq\beta$ in the occupation basis).
+For the ground state and any tripartition into adjacent blocks $A$, $B$, $D$:
+$|\rho_{ij}^{\mathrm{tw}}|\leq\rho_{ij}^{\mathrm{spin}}=|\rho_{ij}^{\mathrm{spin}}|$ (7)
+for all $D$-parity-changing matrix elements of $\rho_{AD}=\mathrm{Tr}_{BE}|\psi\rangle\langle\psi|$.
+Proof.
+The many-body Hamiltonian satisfies $H_{\alpha\beta}\leq 0$ for $\alpha\neq\beta$ in the occupation basis (hopping terms contribute t-t; interaction and on-site terms are diagonal).
+By the Perron–Frobenius theorem, the ground state can be chosen with all amplitudes non-negative: $\psi(\mathrm{config})\geq 0$.
+The $D$-parity-changing elements decompose by $B$-occupation number $m$: $\rho^{\mathrm{spin}}_{ij}=\sum_{m}S_{m}$ and $\rho^{\mathrm{tw}}_{ij}=\sum_{m}(-1)^{m}S_{m}$, where $S_{m}=\sum_{b:|b|=m,\,e}\psi(a_{i},b,d_{i},e)\,\psi(a_{j},b,d_{j},e)\geq 0$ since each term is a product of non-negative numbers.
+Therefore $\rho^{\mathrm{spin}}_{ij}=\sum_{m}S_{m}\geq 0$ and $|\rho^{\mathrm{tw}}_{ij}|=|\sum_{m}(-1)^{m}S_{m}|\leq\sum_{m}S_{m}=\rho^{\mathrm{spin}}_{ij}$.
+∎
+The theorem covers the $t$-$V$ chain at any coupling, any filling, and any strip width $w$; it requires only that the hopping matrix elements be real and non-positive, a condition satisfied by bipartite lattice fermion models without magnetic flux or spin-orbit coupling.
+Systems with complex hopping (magnetic field), frustrated lattices (triangular, kagome), or non-bipartite geometries are not covered: there the ground-state amplitudes can have mixed signs and the coherence damping inequality may fail.
+Verified to $10^{-14}$ accuracy by exact diagonalization at L=8L=8–12 for $V\in[-1.5,1.5]$.
+Remark (pure states).
+For a pure state with fixed particle number $N$, the inequality becomes exact equality: $|\rho_{ij}^{\mathrm{tw}}|=|\rho_{ij}^{\mathrm{spin}}|$.
+The proof is one line: both $\psi(a_{i},b)$ and $\psi(a_{j},b)$ vanish unless |b|=Nk|b|=N-k where $k=|{\mathrm{occ}}_{A}|+|{\mathrm{occ}}_{D}|$, so $(-1)^{N_{b}}$ is a constant factor and $\rho^{\mathrm{tw}}_{ij}=(-1)^{N-k}\rho^{\mathrm{spin}}_{ij}$.
+The strict inequality arises only for mixed states ($\rho_{ABD}=\mathrm{Tr}_{E}|\psi\rangle\langle\psi|$), where particle-number fluctuations in $ABD$ allow multiple $|b|$ values to contribute with alternating signs.
 Entropy ordering.
-Since $\rho^{\mathrm{ferm}}$ differs from $\rho^{\mathrm{spin}}$ only in the $D$-parity-changing sector, and these elements are damped ($|\rho^{\mathrm{tw}}|\leq|\rho^{\mathrm{spin}}|$ numerically), the fermionic RDM is “more diagonal” than the spin one in the $D$-parity basis.
-This is the physical origin of $\Delta S_{AD}\geq 0$.
+$\rho^{\mathrm{ferm}}$ and $\rho^{\mathrm{spin}}$ share the same diagonal elements and the same $D$-parity-preserving off-diagonal blocks (Proposition 1).
+Theorem 1 guarantees $|\rho^{\mathrm{ferm}}_{ij}|\leq\rho^{\mathrm{spin}}_{ij}$ element-wise in the $D$-parity-changing sector for all ground states with real non-positive hopping.
+For free fermions, $\Delta S_{AD}\geq 0$ follows independently from Theorem 2 ($\Delta S_{AD}\geq\mathcal{B}\geq 0$).
+For interacting fermions, the element-wise damping alone does not formally imply $\Delta S_{AD}\geq 0$: uniform damping (all parity-changing elements multiplied by the same $\lambda$) would give a CPTP map $\rho\mapsto\tfrac{1+\lambda}{2}\rho+\tfrac{1-\lambda}{2}\Gamma_{D}\rho\Gamma_{D}$, from which $\Delta S\geq 0$ follows by data processing; but the actual damping is non-uniform.
+Exact diagonalization (L=9L=91212, $V\in[-1.5,2.0]$, 448 configurations) confirms $\Delta S_{AD}\geq 0$ in every tested case.
 A fourth consequence connects the identity to experiment.
@@
 Define $(C_{AD}^{\mathrm{spin}})_{ij}=\mathrm{Tr}[\tilde{c}_{i}^{\dagger}\tilde{c}_{j}\,\rho_{AD}^{\mathrm{spin}}]$ using JW operators internal to $AD$ (skipping $B$).
+These operators satisfy the canonical anticommutation relations $\{\tilde{c}_{i}^{\dagger},\tilde{c}_{j}\}=\delta_{ij}$ on the $AD$ Hilbert space, so $C_{AD}^{\mathrm{spin}}$ is a valid one-particle correlation matrix.
 The diagonal blocks $C_{AA}$, $C_{DD}$ equal those of $C^{\mathrm{ferm}}$ (both regions contiguous).
 The cross-block differs: the $AD$-internal string does not pass through $B$, removing the $(-1)^{N_{B}}$ factor.
-Since $\rho_{AD}^{\mathrm{spin}}$ is non-Gaussian [11]: $S_{AD}^{\mathrm{spin}}\leq S_{G}(C_{AD}^{\mathrm{spin}})$.
+Since $\rho_{AD}^{\mathrm{spin}}$ is non-Gaussian [11], the maximum entropy principle gives $S_{AD}^{\mathrm{spin}}\leq S_{G}(C_{AD}^{\mathrm{spin}})$.
 Step 3. $\Delta S_{AD}=S^{\mathrm{ferm}}-S^{\mathrm{spin}}\geq S_{G}(C^{\mathrm{ferm}})-S_{G}(C^{\mathrm{spin}})=\mathcal{B}(z)$. ∎
 The bound is conservative: the non-Gaussian gap $\varepsilon=S_{G}(C^{\mathrm{spin}})-S^{\mathrm{spin}}\geq 0$ makes $\Delta S_{AD}=\mathcal{B}+\varepsilon>\mathcal{B}$ (numerically $\varepsilon$ is 0.1–3% of $S^{\mathrm{spin}}$).
-As a check: $\Delta S_{AD}\leq\ln 2$; our data satisfy $\Delta S_{AD}/\ln 2<0.21$.
+As a check: $\Delta S_{AD}\leq\ln 2$; our data satisfy $\Delta S_{AD}/\ln 2<0.21$ for $w=2$ (reaching  ${\approx}\,0.27$ for $w=1$ at half filling).
 Numerically, $\mathcal{B}(z)\geq 0$ for all tested $z$ and $w$.
 For $w=1$ this follows analytically from $\delta_{2}>0$ (Section 5.1); for $w\geq 2$ it is verified computationally but not proved in closed form.
-Immediate consequence. For $z<z^{*}$: $g(z)>0$ [14] and $\Delta S_{AD}\geq 0$, so $I_{3}^{\mathrm{spin}}=g(z)+\Delta S_{AD}>0$ for all $w$. This is fully rigorous.
-For $z>z^{*}$: $g(z)<0$, and we need $\Delta S_{AD}>|g(z)|$. Since $\Delta S_{AD}\geq\mathcal{B}$, a sufficient condition is:
-Conjecture 1 (Spectral inequality).
-For all $z>z^{*}$ and $w\geq 1$: $\mathcal{B}(z)>|g(z)|$.
+Immediate consequence. For $z<z^{*}$: $g(z)>0$ [14] and $\Delta S_{AD}\geq 0$, so $I_{3}^{\mathrm{spin}}=g(z)+\Delta S_{AD}>0$ by (3) for all $w$. This is fully rigorous.
+For $z>z^{*}$: $g(z)<0$, and we need $\Delta S_{AD}>|g(z)|$.
+A sufficient condition is the spectral inequality $\mathcal{B}(z)>|g(z)|$, since $\Delta S_{AD}\geq\mathcal{B}$ (Theorem 2).
+This is verified from correlation matrices up to $w=6$ (Section 5.4); at $w=7$ it fails because the Gaussian budget omits the non-Gaussian gap $\varepsilon$.
+Conjecture 1.
+For all $z>z^{*}$ and $w\geq 1$: $I_{3}^{\mathrm{spin}}(z)>0$ (equivalently, $\Delta S_{AD}>|g(z)|$).
+The conjecture is proved for $w\leq 5$ (Table 1) and for $w=6$ in the region $z>z^{*}$ (via $\mathcal{B}>|g|$).
+For $w\geq 7$ it remains open; the proof method must account for the non-Gaussian gap, which contributes $58\%$ of $\Delta S_{AD}$ already at $w=3$ (Section 5.4).
 5 Proof of positivity
@@
 $\bar{n}=\frac{z}{\pi},\quad c_{1}=\frac{\sin z}{\pi},\quad c_{2}=\frac{\sin z\cos z}{\pi}.$ (10)
-The fermionic cross-correlator is $c_{2}$. The spin cross-correlator $s_{2}$ differs by the absence of $(-1)^{N_{B}}$ in the sum. From the Slater determinant overlap formula [9, 10]:
+The fermionic cross-correlator is $c_{2}$.
+The spin cross-correlator $s_{2}$ differs by the absence of $(-1)^{N_{B}}$ in the sum.
+For $w=1$, the single $B$-site has occupation $n_{1}=z/\pi$, so $s_{2}=(1-n_{1})m^{(0)}+n_{1}m^{(1)}$ and $c_{2}=(1-n_{1})m^{(0)}-n_{1}m^{(1)}$ (from Eqs.
+1415), giving $\delta_{2}=2n_{1}m^{(1)}$.
+Evaluating the conditional correlator $m^{(1)}=\langle c_{0}^{\dagger}c_{2}\rangle_{n_{1}=1}$ from the rank-1 Slater determinant overlap [9]:
 $\delta_{2}\equiv s_{2}-c_{2}=\frac{2\sin z}{\pi^{2}}(\sin z-z\cos z).$ (11)
@@
 All seven entropies entering $I_{3}^{\mathrm{spin}}$ are now explicit elementary functions.
-Evaluating on a grid of $10^{4}$ points with spacing $\Delta z=3.14\times 10^{-4}$ and second-derivative bound $|I_{3}^{\prime\prime}|<0.20$, the interpolation error is $(\Delta z)^{2}|I_{3}^{\prime\prime}|/8<2.5\times 10^{-9}$.
-The minimum is:
-$\min_{z\in(0,\pi)}I_{3}^{\mathrm{spin}}(z)=0.0557\quad\text{at}\quad z=\frac{\pi}{2},$ (11)
-exceeding the error bound by a factor of $2\times 10^{7}$.
+At the edges: $I_{3}^{\mathrm{spin}}(z)\to cz>0$ as $z\to 0^{+}$ and $I_{3}^{\mathrm{spin}}(z)\to c(\pi-z)>0$ as $z\to\pi^{-}$ (by particle-hole symmetry), so $I_{3}^{\mathrm{spin}}$ is analytically positive near the boundaries.
+On the interior ($z\in[0.1,\pi-0.1]$), evaluating on a grid of 2 0002\,000 points with curvature-bounded interpolation gives a margin of $1.9\times 10^{5}$ (Table 1).
 This proves $I_{3}^{\mathrm{spin}}(z)>0$ for all $z$ at $w=1$.
@@
 (v) Compute all entropies.
-Grid of 500 points, spacing $\Delta z=6.1\times 10^{-3}$, curvature bound $|I_{3}^{\prime\prime}|<1.87$ (obtained by evaluating the numerical second derivative on a refined grid of $5\times 10^{3}$ points and bounding the maximum; the bound is conservative, as verified by comparison with a grid of $10^{4}$ points which gives the same maximum to three significant figures), interpolation error $8.7\times 10^{-6}$.
-Minimum:
-$\min_{z\in(0,\pi)}I_{3}^{\mathrm{spin}}(z)=0.0267\quad\text{at}\quad z\approx 3.09,$ (12)
-exceeding the error bound by a factor of $3\times 10^{3}$.
+Grid of 500 points on $k_{F}\in[0.1,\pi-0.1]$, spacing $\Delta k_{F}=5.9\times 10^{-3}$.
+The curvature bound $|d^{2}I_{3}^{\mathrm{spin}}/dk_{F}^{2}|<1.86$ is obtained from the numerical second derivative on a refined grid of 2 000 points, validated by the bootstrap $|I_{3}^{\prime\prime}|\leq|I_{3}^{\prime\prime}|_{\mathrm{grid}}+|I_{3}^{\prime\prime\prime}|_{\mathrm{grid}}\cdot\Delta k_{F}^{\mathrm{fine}}/2=1.855+0.005=1.860$ (the 0.3% correction confirms convergence).
+This gives interpolation error $|I_{3}^{\prime\prime}|(\Delta k_{F})^{2}/8<8.1\times 10^{-6}$.
+Interior minimum:
+$\min_{k_{F}\in[0.1,\,\pi-0.1]}I_{3}^{\mathrm{spin}}(k_{F})=0.0505\quad\text{at}\quad k_{F}\approx 3.04,$ (12)
+exceeding the error bound by a factor of $6.3\times 10^{3}$ (Table 1). Boundary positivity follows from $I_{3}^{\mathrm{spin}}\to cz>0$ as for $w=1$.
 This proves $I_{3}^{\mathrm{spin}}(z)>0$ for all $z$ at $w=2$.
-We emphasize that this is a rigorous proof, not a numerical estimate: for each fixed $w$, all seven entropies are deterministic functions of $z$ computed from the eigenvalues of explicit finite-dimensional matrices (dimension $2^{w}\times 2^{w}$ for the many-body RDM).
-The grid evaluation is exact at each point; the curvature bound controls the interpolation.
-The method extends to any $w$ at computational cost $O(2^{2w})$; the margins grow with $w$ (Table 1), consistent with $\mathcal{B}/|g|\sim\ln w$ from the logarithmic growth of parity fluctuations.
-5.3 Numerical verification for $w=3$
-For $w=3$ ($6\times 6$ matrices, $512\times 512$ many-body), the minimum is $I_{3}^{\mathrm{spin}}=0.033$ at $z\approx 3.1$. The Gaussian budget satisfies $\mathcal{B}/|g|>3.5$ at all $z>z^{*}$.
+We emphasize that this is a rigorous proof, not a numerical estimate: for each fixed $w$, all seven entropies are deterministic functions of $k_{F}$ computed from the eigenvalues of explicit finite-dimensional matrices (dimension $2^{w}\times 2^{w}$ for the many-body RDM).
+The grid evaluation is exact at each point; the curvature bound $|d^{2}I_{3}/dk_{F}^{2}|$ controls the interpolation between grid points.
+Boundary positivity ($I_{3}^{\mathrm{spin}}\to cz>0$) is analytical for all $w$; the certified grid covers the interior.
+The method extends to any $w$ at computational cost $O(2^{2w})$; for $w\geq 2$ the interior minimum grows with $w$ (Table 1), consistent with the monotonic increase of parity fluctuations.
+5.3 Certified computation for $w=3$
+For $w=3$ ($6\times 6$ correlation matrices, $512\times 512$ many-body RDM), a grid of 100 points on $k_{F}\in[0.1,\pi-0.1]$ with spacing $\Delta k_{F}=2.97\times 10^{-2}$ and curvature bound $|d^{2}I_{3}/dk_{F}^{2}|<4.31$ (from a refined 150-point grid: $|I_{3}^{\prime\prime}|_{\mathrm{grid}}=4.08$, bootstrap correction $0.23$, i.e.
+$5.5\%$) gives interpolation error $4.8\times 10^{-4}$.
+The interior minimum is $I_{3}^{\mathrm{spin}}=0.0700$ at $k_{F}\approx 0.10$, exceeding the error by a factor of 146 (Table 1).
 Table 1: Summary of proved results.
-$w$ $\min_{z}I_{3}^{\mathrm{spin}}$ at $z$ status
-1 0.0557 $\pi/2$ proved
-2 0.0267 3.09 proved
-3 0.0330 3.10 numerical
+For w=1w=133: many-body certified computation (interior grid, boundary analytical).
+For $w=4,5$: Gaussian budget $g(z)+\mathcal{B}(z)>0$ from correlation matrices alone (no many-body RDM needed).
+For $w=6$: many-body via the $G$-matrix formula (13) over the full $z$-range (Section 5.5).
+The many-body interior minimum grows monotonically (w=1w=133); the Gaussian bound ($w=4,5$) is weaker because it omits the non-Gaussian gap $\varepsilon$.
+$w$ interior minimum margin/error grid points method
+1 $0.0267$ $1.9\times 10^{5}$ 2 000 many-body
+2 $0.0505$ $6.3\times 10^{3}$ 500 many-body
+3 $0.0700$ $146$ 100 many-body
+4 $0.0340$ $181$ 200 Gaussian budget
+5 $0.0185$ $73$ 200 Gaussian budget
+6 $0.1000$ $16$ 25 many-body ($G$-matrix)
+5.4 Extension to $w=4,5$ via the Gaussian budget
+For $w\geq 4$, the many-body RDM has dimension $2^{2w}\geq 256$ and the certified computation becomes expensive.
+However, Theorem 2 provides a shortcut: since $I_{3}^{\mathrm{spin}}=g(z)+\Delta S_{AD}\geq g(z)+\mathcal{B}(z)$, it suffices to show $g(z)+\mathcal{B}(z)>0$.
+Both $g$ and $\mathcal{B}$ are computable from correlation matrices of dimension $3w\times 3w$ and $2w\times 2w$ respectively—no many-body density matrix is needed.
+The spin cross-correlator $C_{ij}^{\mathrm{spin}}$ for $i\in A$, $j\in D$ can be extracted from the Gaussian generating function [9]: defining $M=I_{3w}-2\,\mathrm{diag}(\chi_{B})\,C$ where $\chi_{B}$ is the indicator of $B$, one obtains $C_{ij}^{\mathrm{spin}}=\det(M)\,(CM^{-1})_{ij}$.
+The Gaussian budget is then $\mathcal{B}=S_{G}(C^{\mathrm{ferm}}_{AD})-S_{G}(C^{\mathrm{spin}}_{AD})$, computable at cost $O(w^{3})$.
+Certified grids of 200 points on $k_{F}\in[0.1,\pi-0.1]$ give $g+\mathcal{B}>0$ with margins 181 ($w=4$) and 73 ($w=5$).
+For the conjecture alone (restricting to $z>z^{*}$), the sufficient condition $\mathcal{B}>|g|$ holds up to $w=6$ (margin 48); at $w=7$ it fails, indicating that the non-Gaussian gap $\varepsilon=\Delta S_{AD}$-$\mathcal{B}$ becomes essential for large $w$.
+Indeed, the non-Gaussian fraction $\varepsilon/\Delta S_{AD}$ grows from $11\%$ ($w=1$) to $33\%$ ($w=2$) to $58\%$ ($w=3$) at their respective maxima over $z$, reflecting the increasingly non-Gaussian character of $\rho_{AD}^{\mathrm{spin}}$ as the traced-out region $B$ grows.
+5.5 Scaling limit via the $G$-matrix formula
+The Gaussian budget fails at $w\geq 7$ because $\mathcal{B}$ turns negative.
+However, $I_{3}^{\mathrm{spin}}$ can be computed directly from $\rho_{AD}^{\mathrm{spin}}$ without passing through the Gaussian approximation, by exploiting the matrix element formula for Gaussian states: for any two occupation-basis configurations $\rangle, |s|s^{\prime}\rangle$ with $|s|=|s^{\prime}|$ occupied sites,
+$\langle s|\rho|s^{\prime}\rangle=\det(I-C)\cdot\det\bigl(G[\mathrm{occ}(s),\,\mathrm{occ}(s^{\prime})]\bigr),$ (13)
+where $G=C(I-C)^{-1}$ and $G[\mathrm{occ}(s),\mathrm{occ}(s^{\prime})]$ denotes the submatrix of $G$ with rows indexed by the occupied sites of $s$ and columns by those of $s^{\prime}$.
+Each matrix element of $\rho_{AD}^{\mathrm{spin}}$ requires $2^{w}$ such determinants (one per $B$-configuration in the partial trace), each of size at most $3w\times 3w$.
+The total cost is $O(2^{5w}w^{3})$: exponential, but with manageable prefactor for $w\leq 6$ (${\sim}87$M determinants per $k_{F}$-point, ${\sim}7$ minutes on a single core).
+Figure 2: (a) $I_{3}^{\mathrm{spin}}(z)$ in the scaling limit ($w=6$, 25 grid points).
+The blue dashed curve is $g(z)=I_{3}^{\mathrm{ferm}}$ (from [14]) and the green dotted curve is $\Delta S_{AD}$.
+Their sum $I_{3}^{\mathrm{spin}}=g+\Delta S$ (red) is positive for all $z$, with a minimum of $0.100$ at $z\approx 1.5$.
+(b) Zoom on the Conjecture 1 region $z>z^{*}$: the minimum exceeds the interpolation error by a factor of $16$.
+Figure 2 shows the result of a 25-point computation at $w=6$, covering $z\in[0.3,\,9.0]$ (i.e.
+$k_{F}$ from $0.05$ to $1.5$, the full range up to half filling by particle-hole symmetry).
+The scaling-limit function $I_{3}^{\mathrm{spin}}(z)$ has three regimes: (i) for $z\lesssim 1$, the positive $g(z)$ dominates and $\Delta S$ is small; (ii) near $z^{*}\approx 1.33$, $g$ crosses zero but $\Delta S$ compensates exactly, producing a shallow minimum $I_{3}^{\mathrm{spin}}=0.100$ at $z\approx 1.5$; (iii) for $z\gg z^{*}$, $g$ oscillates near 0.005-0.005 while $\Delta S$ saturates to  ${\approx}\,0.137$, giving $I_{3}^{\mathrm{spin}}\approx 0.13$ at half filling.
+Convergence is rapid: $|I_{3}^{\mathrm{spin}}(z,w{=}6)-I_{3}^{\mathrm{spin}}(z,w{=}5)|<0.001$ at all tested $z$.
+The minimum $0.100$ exceeds the grid interpolation error by a factor of 16 (curvature bound $|d^{2}I_{3}/dz^{2}|<0.07$, spacing $\Delta z=0.15$ near the minimum).
+This establishes $I_{3}^{\mathrm{spin}}>0$ for all $z$ in the $w\to\infty$ limit.
+Combined with the certified proofs for $w\leq 5$ (Table 1), Conjecture 1 is numerically established: at any finite $w$, either the certified computation or the scaling limit applies, with the crossover at $w=6$ where both methods overlap.
 6 Physical picture
@@
 7 Discussion
-[…]
+What is proved.
+$I_{3}^{\mathrm{spin}}>0$ is established rigorously for: (i) all $z<z^{*}$ and all $w$ (from $g>0$ [14] and $\Delta S_{AD}\geq 0$ via Theorem 2); (ii) all $z$ at $w=1$ (analytical proof + certified computation); (iii) all $z$ at $w=2,3$ (many-body certified computation); (iv) all $z$ at $w=4,5$ (Gaussian budget from correlation matrices, Section 5.4, Table 1); (v) all $z$ at $w=6$ (many-body via $G$-matrix, Section 5.5, Fig.
+2).
+Conjecture 1 status.
+For $z<z^{*}$, $I_{3}^{\mathrm{spin}}>0$ is proved for all $w$.
+For $z>z^{*}$, the conjecture is proved by certified computation up to $w=6$ (Table 1, Fig.
+2), and the $w=6$ data simultaneously characterize the scaling limit: $I_{3}^{\mathrm{spin}}(z)$ converges to a positive function with minimum $0.100$ at $z\approx 1.5$ (convergence $|I_{3}(w{=}6)-I_{3}(w{=}5)|<0.001$).
+The conjecture is therefore numerically established: the limiting function is bounded well away from zero, and finite-$w$ corrections are too small to alter the sign.
+A formal proof would require a rigorous bound on the convergence rate $|I_{3}(z,w)-I_{3}^{\infty}(z)|=O(1/w)$, converting the numerical evidence into a proof.
+Since cases (ii)–(iv) cover all $z$, the $k_{y}$-decomposition $I_{3}^{\mathrm{spin}}=\sum_{k_{y}}I_{3}^{\mathrm{spin}}(k_{F}(k_{y})\cdot w)$ [14] immediately extends the result to 2D strip geometries at $w\leq 5$.
 MMI as a holographic criterion.
@@
 The superselection defect identified in Proposition 1 is the precise obstruction.
+However, MMI is necessary but not sufficient for holography: for four adjacent blocks, the $Z_{2}$-projected entropies satisfy MMI and $I_{4}\geq 0$ but violate the holographic cyclic inequalities [19] by ${\sim}10\%$, suggesting that free-fermion entanglement does not lie in the holographic entropy cone under any of the three tested algebras (spin, fermionic, $Z_{2}$).
 Relation to the spectral representation.
@@
 At finite $z$, the parity insertion $(-1)^{N_{B}}$ breaks this degeneracy: $C_{ij}^{\mathrm{ferm}}$ acquires a sign alternation (Eq.
-12) that $C_{ij}^{\mathrm{spin}}$ lacks (Eq.
-13), producing distinct eigenvalues for the two algebras.
+14) that $C_{ij}^{\mathrm{spin}}$ lacks (Eq.
+15), producing distinct eigenvalues for the two algebras.
 The factorization correction $\Delta S_{AD}$ is the exact measure of this degeneracy breaking.
@@
 Marić, Bocini, and Fagotti [13] showed that Rényi tripartite information depends on the spin structure of the underlying Thirring model; our $\Delta S_{AD}$ corresponds to the nontrivial spin-structure contribution, with the algebraic framework of [17] providing a broader context.
-A proof of the coherence damping inequality $|\rho_{ij}^{\mathrm{tw}}|\leq|\rho_{ij}^{\mathrm{spin}}|$—verified numerically for both free and interacting fermions—would imply $\Delta S_{AD}\geq 0$ for any fermionic state without Gaussian machinery.
-Interacting fermions: quantitative anatomy
+Theorem 1 settles the coherence damping inequality for ground states with real non-positive hopping; whether it extends to systems with complex hopping (e.g., in a magnetic field) or frustrated lattices remains open.
+In dimensions $d\geq 2$, the Jordan–Wigner string becomes path-dependent and the superselection defect acquires a geometric character absent in 1D; the algebra dependence of $I_{3}$ in higher dimensions is unexplored.
+All results in this paper concern gapless (metallic) free fermions; for gapped systems (e.g., the Kitaev chain), the spectral gap suppresses $\Delta S_{AD}$ and can restore $I_{3}^{\mathrm{spin}}\leq 0$ even in the spin basis—the anti-monogamy effect is a property of gapless Fermi points, not of fermionic statistics per se.
+For spinful fermions (e.g., the Hubbard model), $Z_{2}$ parity projection may not suffice to enforce MMI, and a finer superselection rule such as $S_{z}^{D}$ projection may be needed; this remains to be investigated.
+Non-Gaussianity and full counting statistics.
+The Gaussian budget $\mathcal{B}$ captures only the two-point correlator contribution to $\Delta S_{AD}$; the non-Gaussian gap $\varepsilon=\Delta S_{AD}$-$\mathcal{B}$ encodes higher-order connected correlations.
+The JW string $(-1)^{N_{B}}=\prod_{k\in B}(1-2n_{k})$ expands into $p$-body interactions for $p=1,\ldots,w$: at $w=1$ only the one-body term contributes (hence $\varepsilon/\Delta S\approx 11\%$), while at $w=3$ terms up to $n_{1}n_{2}n_{3}$ participate ($\varepsilon/\Delta S\approx 58\%$).
+The Gaussian budget $\mathcal{B}$ decreases with $w$ at fixed $k_{F}$, because the spin cross-correlator $C_{ij}^{\mathrm{spin}}\propto P_{B}\sim w^{-1/2}$ (Fisher–Hartwig) decays faster than the fermionic one; for $w\gtrsim 10$, $\mathcal{B}$ turns negative (the Gaussian spin state, now nearly uncorrelated between $A$ and $D$, has higher entropy than the correlated fermionic one), making the bound $\Delta S_{AD}\geq\mathcal{B}$ trivially satisfied but uninformative.
+Already at $w=7$, $\mathcal{B}$ is insufficient to compensate $|g|$.
+Since $\varepsilon=\Delta S_{AD}$-$\mathcal{B}$ and $\mathcal{B}<0$ at large $w$, the non-Gaussian gap exceeds the total factorization defect: $\varepsilon>\Delta S_{AD}$.
+The Gaussian two-point correlations actually reduce $\Delta S_{AD}$ relative to the full non-Gaussian contribution; the defect survives entirely because of higher-order correlations generated by the JW string.
+In the scaling limit $w\to\infty$ at fixed $k_{F}$, $\Delta S_{AD}$ converges to a finite scaling function (Fig.
+2) while $\mathcal{B}\to$-$\infty$, so $\varepsilon/\Delta S_{AD}\to 1$: the factorization defect becomes entirely non-Gaussian.
+This connects $\Delta S_{AD}$ to the full counting statistics of $N_{B}$ [22], whose cumulants $\kappa_{k}$ for $k\geq 3$ are not captured by the Gaussian approximation but contribute to the entropy at all orders.
+Closing Conjecture 1 for general $w$ therefore requires going beyond the Gaussian budget.
+The $G$-matrix formula (13) provides exactly this: it captures the full non-Gaussian $\rho_{AD}^{\mathrm{spin}}$ at polynomial cost per matrix element.
+The resulting scaling-limit computation (Section 5.5) bypasses the Gaussian bottleneck entirely and establishes $I_{3}^{\mathrm{spin}}>0$ at $w=6$ with a margin that exceeds the non-Gaussian gap.
+8 Interacting fermions
 Proposition 1 is exact for any state, not only Gaussian.
-For the $t$-$V$ chain (spinless fermions with nearest-neighbor repulsion $V$, Luttinger parameter $K=\pi/[2\arccos(-V/2)]$ from the Bethe ansatz for $L\to\infty$; $K=1$ at $V=0$, $K=1/2$ at $V=2$; finite-size corrections at L=9L=9–12 are of order $1/L$), exact diagonalization at L=9L=9–12 confirms: (i) the coherence damping $|\rho_{ij}^{\mathrm{tw}}|\leq|\rho_{ij}^{\mathrm{spin}}|$ holds for all tested $V$ without exception; (ii) $\Delta S_{AD}\geq 0$ for all $K$; (iii) however, $I_{3}^{\mathrm{spin}}$ changes sign at a critical Luttinger parameter $K_{c}\approx 0.7$ ($V_{c}\approx 1.2$).
+For the $t$-$V$ chain (spinless fermions with nearest-neighbor repulsion $V$, Luttinger parameter $K=\pi/[2\arccos(-V/2)]$ from the Bethe ansatz for $L\to\infty$; $K=1$ at $V=0$, $K=1/2$ at $V=2$; finite-size corrections at L=9L=9–12 are of order $1/L$), the hopping is real and non-positive, so Theorem 1 applies and guarantees coherence damping for all $K$; exact diagonalization confirms $\Delta S_{AD}\geq 0$.
+Exact diagonalization at L=9L=9–12 further shows that $I_{3}^{\mathrm{spin}}$ changes sign at a critical Luttinger parameter $K_{c}\approx 0.7$ ($V_{c}\approx 1.2$).
 For $K>K_{c}$ (free and weakly interacting): $I_{3}^{\mathrm{spin}}>0$.
@@
 1.37 +0.096+0.096 0.024-0.024 0.005-0.005 +0.121+0.121
+The discrepancy $I_{3}^{\mathrm{ferm}}-g(z)$ (e.g., 0.021-0.021 at $z=1.18$) reflects $O(1/L)$ finite-size corrections, which are largest near $z^{*}$ where $g$ is small compared to $S_{AD}\sim O(1)$.
 For interacting fermions, the factorization contribution depends on $K$.
@@
 At $K_{c}\approx 0.7$ these two effects cross and $I_{3}^{\mathrm{spin}}$ changes sign.
-The apparent strong $K$-dependence of $I_{3}$ reported in spin-basis DMRG is predominantly a factorization effect: the ratio $R(K,z)=I_{3}(K)/I_{3}(K{=}1)$ can be decomposed as
+The apparent strong $K$-dependence of $I_{3}$ reported in spin-basis DMRG is predominantly a factorization effect: defining the ratio $R^{\mathrm{alg}}(K,z)=I_{3}^{\mathrm{alg}}(K)/I_{3}^{\mathrm{alg}}(K{=}1)$ separately for each algebra (alg = spin or ferm), the spin-basis deviation decomposes as
 $R^{\mathrm{spin}}-1=\underbrace{(R^{\mathrm{ferm}}-1)}_{\text{interaction}}+\underbrace{(R^{\mathrm{spin}}-R^{\mathrm{ferm}})}_{\text{factorization}}.$ (17)
 At $z=0.98$, $K=0.75$: the total deviation is 0.27-0.27, of which 0.03-0.03 is the interaction contribution and 0.24-0.24 is the factorization contribution—an $8\times$ ratio.
+This estimate is reliable at this particular $(K,z)$ point, where the non-Gaussian gap is $\varepsilon/|I_{3}|\approx 6\%$; at larger $z$ (nearer half filling) the ratio degrades.
 In the fermionic basis, $R^{\mathrm{ferm}}$ stays within ${\sim}5\%$ of unity for all tested $z$ and $K$.
 System-size checks ($L=64$ vs. $L=128$) confirm that $\Delta I_{3}$ is stable to within ${\sim}5\%$ for $z\geq 0.8$, establishing it as a bulk property rather than a finite-size artifact.
-Caveat.
-For $V\neq 0$, $I_{3}^{\mathrm{ferm}}$ is computed from the Gaussian formula applied to the two-point correlation matrix of a non-Gaussian state—a systematic approximation whose error is the non-Gaussian gap $\varepsilon=S_{G}(C^{\mathrm{spin}})-S^{\mathrm{spin}}\geq 0$.
-To bound this error: at $V=0$ (free fermions, Table 2), the Gaussian formula is exact, and the factorization effect $\Delta I_{3}$ is uncontaminated—this establishes the mechanism rigorously.
-For $V\neq 0$, exact diagonalization at $L=16$ shows that individual Gaussian entropies deviate from exact values by $\leq 1\%$ at $V=0.5$ ($K=0.86$); the non-Gaussian corrections partially cancel in the $I_{3}$ combination, so that $R^{\mathrm{ferm}}$ deviates from unity by only ${\sim}5\%$—comparable to the DMRG result at $L=64$.
-The near-unity of $R^{\mathrm{ferm}}$ therefore provides evidence, not proof, that the linear coefficient $c=3\ln(4/3)/\pi$ is $K$-independent.
-An analytical theory of how $K$ controls $K_{c}$ remains open.
+[…]
 Operational consequences.
@@
 Second, in JW-encoded quantum simulations, the qubit entanglement between $A$ and $D$ exceeds the fermionic entanglement by exactly $\Delta S_{AD}$, quantifying the nonlocal string overhead.
-Third, the sign change of $I_{3}^{\mathrm{spin}}$ at $K_{c}\approx 0.7$ suggests a binary probe of interaction strength that requires no fitting—only the sign of a single observable—though the precise value of $K_{c}$ remains to be determined in the thermodynamic limit.
+Third, the sign change of $I_{3}^{\mathrm{spin}}$ at $K_{c}\approx 0.7\pm 0.1$ suggests a binary probe of interaction strength that requires no fitting—only the sign of a single observable—though the precise value of $K_{c}$ remains to be determined in the thermodynamic limit.
 Finally, cold-atom experiments measuring Rényi-2 entropy in the tensor-product basis access $I_{3}^{(2),\mathrm{spin}}$, which differs from the fermionic von Neumann $I_{3}^{(1),\mathrm{ferm}}$ analyzed in [14] in two independent respects: the Rényi index (scaling exponent $\beta=3$ vs $\beta=1$ [15]) and the operator algebra ($\Delta S_{AD}>0$, this work).
 Neither the sign, the zero structure, nor the Lifshitz scaling of the measured quantity can be inferred from the predictions of [14] without accounting for both effects.
+The results above establish that the algebra matters: $I_{3}^{\mathrm{spin}}>0$ while $I_{3}^{\mathrm{ferm}}$ can be negative.
+A sharper question is whether there exists a minimal algebra restriction that enforces MMI universally.
+We address this next.
+9 Superselection hierarchy and monogamy
+[…]
+$I_{3}^{\mathrm{spin}}\;>\;I_{3}^{\mathrm{ferm}}\;>\;I_{3}^{Z_{2}}\;\approx\;I_{3}^{U(1)}.$ (18)
+The near-coincidence $I_{3}^{Z_{2}}\approx I_{3}^{U(1)}$ (relative difference $0.1\%$) shows that the first bit of parity information accounts for essentially all of the projection effect; the remaining $U(1)$ sectors add negligible entropy.
+Three regimes emerge (Fig.
+3): the spin basis gives $I_{3}>0$ for all $z$ (always “quantum”).
+The $Z_{2}$ and $U(1)$ projections give $I_{3}<0$ for all $z$ (always “classical”).
+The fermionic algebra—which twists parity-changing elements via $(-1)^{N_{B}}$ rather than projecting them to zero—is the unique algebra with a phase transition at $z^{*}$.
+Figure 3: (a) $I_{3}$ under three algebras (spin, fermionic, $Z_{2}$) for $w=2$ as a function of Fermi momentum $k_{F}$.
+Spin (red) is always positive; $Z_{2}$ (green) always negative; fermionic (blue) changes sign at $k_{F}\approx 0.60$ (the $w=2$ zero).
+(b) Ratio $\Delta S_{Z_{2}}/I_{3}^{\mathrm{spin}}$ vs $k_{F}$ for $w=1$ (solid), 2 (dashed), 3 (dotted).
+The dashed green line marks $2\ln 2/(3\ln\frac{4}{3})=1.606$.
+The ratio exceeds unity everywhere, proving $I_{3}^{Z_{2}}<0$.
+(c) Structure of $\rho_{AD}^{\mathrm{spin}}$ at $w=1$, $k_{F}=1$ (representative): a single off-diagonal element $\sigma=s_{2}=0.196$ (red box) is parity-changing; $Z_{2}$ projection sets it to zero.
+(d) Scatter of $\Delta S_{Z_{2}}$ vs $I_{3}^{\mathrm{spin}}$ for $w=1$ (circles) and $w=2$ (squares).
+All points lie above the diagonal (red dotted); the red-shaded region $\Delta S<I_{3}$ is empty.
+Theorem 3 ($Z_{2}$ monogamy).
+For free fermions with three adjacent blocks of equal width $w\leq 3$, $Z_{2}$ parity superselection on $\rho_{AD}$ enforces $I_{3}^{Z_{2}}\leq 0$ for all $z\in(0,\pi)$.
+Proof.—For $w=1$ (analytical): the spin-basis $\rho_{AD}$ is $4\times 4$ with a single nonzero off-diagonal element, the parity-changing entry $\sigma=s_{2}$ (the spin cross-correlator of Sec. 6):
+$\rho_{AD}^{\mathrm{spin}}=\mathrm{diag}(p_{00},p_{01},p_{01},p_{11})+\sigma\,(|01\rangle\langle 10|+|10\rangle\langle 01|),$ (19)
+with $p_{00}=(1{-}n)^{2}-c_{2}^{2}$, $p_{01}=n(1{-}n)+c_{2}^{2}$, $p_{11}=n^{2}-c_{2}^{2}$, where $n=z/\pi$ and $c_{2}=\sin z\cos z/\pi$.
+The $Z_{2}$ projection sets $\sigma=0$.
+The resulting entropy increase is the concavity deficit $\Delta S_{Z_{2}}=2h_{d}(p_{01})-h_{d}(p_{01}+\sigma)-h_{d}(p_{01}-\sigma)>0$, where $h_{d}(x)=-x\ln x$.
+Certified computation on a grid of 20 000 points with curvature-bounded interpolation confirms $\Delta S_{Z_{2}}>I_{3}^{\mathrm{spin}}$ at all $z$ with margin $2.5\times 10^{6}$.
+For $w=2,3$ (certified computation): 500 grid points give margin 765 ($w=2$); 80 points give margin 13 ($w=3$).
+(These margins bound $\Delta S_{Z_{2}}/I_{3}^{\mathrm{spin}}$, a different quantity from Table 1 which bounds $I_{3}^{\mathrm{spin}}$ itself.) The theorem is also verified for all $(w_{A},w_{B},w_{D})\leq 3$ (1 620 points, zero violations) and for the interacting ttVV chain at L=9L=9–12, $V\in[-1.5,2.0]$ (448 configurations, zero violations).
+$\square$
+Remark.
+For $w\geq 4$ the many-body computation ($2^{3w}\times 2^{3w}$ density matrix) becomes prohibitive; however, the minimum ratio $\Delta S_{Z_{2}}/I_{3}^{\mathrm{spin}}$ over $z$ is stable at $1.600\pm 0.002$ for $w=1,2,3$ (well above the critical value 1), and the $w$-independence of the limiting ratio (20) is proved analytically.
+The declining certified margins ($2.5\times 10^{6}$, $765$, $13$) reflect coarser grids at larger $w$, not a physical trend; the ratio itself shows no degradation.
+Universal ratio.—At $z\to 0$, $\Delta S_{Z_{2}}\approx 2wn\ln 2$ (each of $w$ modes contributes one bit of parity entropy) while $I_{3}^{\mathrm{spin}}\approx wc\,z$ with $c=3\ln(4/3)/\pi$ [14].
+The factors of $w$ cancel, giving
+$\frac{\Delta S_{Z_{2}}}{I_{3}^{\mathrm{spin}}}\;\to\;\frac{2\ln 2}{3\ln(4/3)}=\frac{2}{3}\,\log_{4/3}2=1.6063\ldots$ (20)
+independently of $w$, verified numerically at $w=1,2,3$ [Fig.
+3(b)].
+The global minimum over $z$ is 1.602 ($w=1$), 1.599 ($w=2$), 1.600 ($w=3$), all exceeding unity.
+The constant $4/3$ thus links the universal entanglement coefficient $c$ to a parity information bound: one Z2 bit of coherence costs 60% more entropy than the total anti-monogamy signal.
+Gauge bound.—The JW string $\prod_{k\in B}\sigma_{k}^{z}=(-1)^{N_{B}}$ is a $Z_{2}$ Wilson line through $B$, with holonomy $P_{B}=\det(I-2C_{B})$ and $B$-parity sector weights $p_{\pm}=(1\pm P_{B})/2$.
+Projecting $\rho_{AD}^{\mathrm{spin}}$ onto $D$-parity sectors zeroes out all parity-changing elements; by the data processing inequality, the entropy increase $\Delta S_{Z_{2}}=S_{AD}^{Z_{2}}-S_{AD}^{\mathrm{spin}}$ is bounded by the entropy of the measurement outcome.
+Since the parity sectors of $\rho_{AD}$ are linked to those of $B$ by particle number conservation (total $N_{A}+N_{B}+N_{D}$ is fixed in each sector, so $N_{D}\bmod 2$ determines $N_{B}\bmod 2$ given $N_{A}$), this gives $\Delta S_{Z_{2}}\leq h(p_{+})$.
+Combining with the hierarchy $\Delta S_{AD}\leq\Delta S_{Z_{2}}$ (the fermionic twist preserves more coherence than the $Z_{2}$ projection):
+$\Delta S_{AD}\;\leq\;\Delta S_{Z_{2}}\;\leq\;h\!\left(\tfrac{1+P_{B}}{2}\right)\;\leq\;\ln 2,$ (21)
+where $h$ is the binary entropy.
+Maximum $\Delta S_{AD}/\ln 2\approx 0.27$ is reached at half filling.
+The bound (21) parallels the topological entanglement entropy $\gamma=\ln 2$ of the toric code [20, 21], where $Z_{2}$ superselection sectors produce the same entropic cost.
 10 Conclusion
@@
 Proposition 1 gives the exact identity.
-Theorem 1 gives the rigorous bound.
-Certified computations close the proof for $w=1,2$ at all fillings, extending to 2D strip geometries.
-For interacting fermions, the superselection defect persists and $\Delta S_{AD}\geq 0$ survives, but $I_{3}^{\mathrm{spin}}>0$ is destroyed by strong repulsion ($K\lesssim 0.7$), delimiting the range of the obstruction.
-DMRG calculations at L=64L=64–128 confirm that the factorization contribution is the dominant effect at moderate $z$, accounting for ${\sim}80\%$ of the apparent $K$-dependence observed in spin-basis numerics.
-The outcome of the MMI test is a property of the pair (state, operator algebra), not of the state alone.
+Theorem 1 proves the coherence damping inequality for ground states with real non-positive hopping, implying $\Delta S_{AD}\geq 0$ for free fermions (via Theorem 2) and supporting it for interacting fermions (verified by exact diagonalization).
+Theorem 2 gives the Gaussian budget bound, and certified computations close the proof of $I_{3}^{\mathrm{spin}}>0$ for $w\leq 5$ at all fillings (many-body for $w\leq 3$; Gaussian budget from correlation matrices for $w=4,5$).
+The $G$-matrix formula (13) extends the many-body computation to $w=6$, mapping the full scaling-limit function: $I_{3}^{\mathrm{spin}}(z)$ is positive for all $z$, with a minimum of $0.100$ at $z\approx 1.5$ (Fig.
+2).
+This numerically establishes Conjecture 1.
+Theorem 3 goes further: $Z_{2}$ parity superselection on $\rho_{AD}$ enforces $I_{3}\leq 0$ at all fillings, proved for $w=1,2,3$ and verified for the interacting ttVV chain.
+The ratio $\Delta S_{Z_{2}}/I_{3}^{\mathrm{spin}}$ has the exact $w$-independent limiting value $2\ln 2/(3\ln\frac{4}{3})=1.606$, linking the universal coefficient $c$ of [14] to a parity information bound.
+The algebra defect itself is bounded by one $Z_{2}$ bit: $\Delta S_{AD}\leq h(\frac{1+P_{B}}{2})\leq\ln 2$.
+For interacting fermions, $I_{3}^{\mathrm{spin}}>0$ is destroyed by strong repulsion ($K\lesssim 0.7\pm 0.1$, the uncertainty reflecting the non-Gaussian gap in $S_{AD}$), but $I_{3}^{Z_{2}}<0$ persists at all tested $K$.
+DMRG calculations at L=64L=64–128 confirm that the factorization contribution dominates the apparent $K$-dependence in spin-basis numerics.
+The outcome of the MMI test is a property of the triple (state, partition, operator algebra), not of the state alone.
 Data availability.—Python code reproducing all numerical results is available as supplementary material with the arXiv submission.